how many different 4-digit numbers can be formed from th digits 2, 3, 4,5,6,7,and 8 if:
a)repetition of digits is not allowed?
b)repetition of digits is allowed?
c)how many of them are odd numbers? refer to (a).
d)how many of them are even numbers? refer to (b).

Answer by sudhanshu_kmr(1152)

there are 4 even and 3 odd numbers....

a)
no. of ways to form 4 digit numbers without repetition = 7P4 = 840


b) no. of ways to form 4 digit numbers with repetition = 7*7*7*7 = 7^4 = 2401


c) we can put odd digit on unit place be 3 ways and can arrange remaining 6
digit on first 3 places by 6P3 ways,
no. of odd numbers ( refer to a ) = 6P3 * 3 = 360



d) for unit place we can put digit by 4 ways and for other 3 places any digit
no. of even numbers when repetition is allowed = 7*7*7*4 =1372